OpenStudy (anonymous):

Chuck and Dana agree to meet for the weekend. Chuck travels 132 miles in the same time that Dana travels 108 miles. If Chuck's rate of travel is 6 mph more that Dana's, and they travel the same length of time, at what speed does Chuck travel?

OpenStudy (anonymous):

x Dana's speed as time is same 108/x=132/(x+6) x=27 Chuck's speed is 33